OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). A. y = four thirdsx + 8 B. y = four thirdsx + 3 C. y = -three fourthsx + 8 D. y = -three fourthsx + 3

OpenStudy (anonymous):

Leave coefficients of x and y the same, then use form: A (x − x0) + B (y − y0) = 0 4 (x + 3) − 3 (y − 4) = 0 4x + 12 − 3y + 12 = 0 −3y = −4x − 24 y = 4/3 x + 8 ------------------------------ Alternate method: Line 4x − 3y has slope 4/3, so parallel line will also have slope 4/3 Parallel line passes through point (−3, 4) Use point−slope form : y − y0 = m (x − x0) y − 4 = 4/3 (x + 3) y = 4/3 x + 4 + 4 y = 4/3 x + 8

OpenStudy (anonymous):

Whoa...

OpenStudy (anonymous):

Thanks!